OpenStudy (anonymous):

An observer (O) is located 900 feet from a building (B). The observer notices a helicopter (H) flying at a 49° angle of elevation from his line of sight. What equation and trigonometric function can be used to solve for the height (h) of the helicopter? Pick another trigonometric function and describe why that function is not appropriate when trying to solve for (h).

OpenStudy (anonymous):

@rishi_84

OpenStudy (anonymous):

since i know the base and want to calculate height so we can use cos cos49 = 900/h h = 900/cos49 = 900/0.65 = 1384.61 we can't use sin because for that we nee hypotaneous that is direct distance between helicopter and observer which is not known

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

Thanks

OpenStudy (mak1204):

@rishi_84 isn't it tan not cos